A manufacturer of light bulbs produces bulbs that last a mean of 950 hours with a standard variation of 120 hours. What is the probability that the mean life-time of a random sample of 10 of these bulbs is less than 900 hours?


1
Expert's answer
2021-03-30T07:53:09-0400

Let "X"  be the random variable which represents the length of the life of a light bulb, in hours. We suppose that "X\sim N(\mu, \sigma^2)." Then "\bar{X}\sim N(\mu, \sigma^2/n)"

Given "\mu=950, \sigma=120, n=10"


"P(X<900)=P(Z<\dfrac{950-900}{120/\sqrt{10}})"

"\approx P(Z<1.317616)\approx0.906184"

The probability that the mean life-time of a random sample of 10 of these bulbs is less than 900 hours is "0.906184."